### v2

# A Layered Protocol Architecture for the Internet of Agents

Abstract.
Large Language Models (LLMs) have demonstrated remarkable performance improvements and the ability to learn domain-specific languages (DSLs), including APIs and tool interfaces. This capability has enabled the creation of AI agents that can perform preliminary computations and act through tool calling, now being standardized via protocols like MCP. However, LLMs face fundamental limitations: their context windows cannot grow indefinitely, constraining their memory and computational capacity. Agent collaboration emerges as essential for solving increasingly complex problems, mirroring how computational systems rely on different types of memory to scale. The ”Internet of Agents” (IoA) represents the communication stack that enables agents to scale by distributing computation across collaborating entities.
Current network architectural stacks (OSI and TCP/IP) were designed for data delivery between hosts and processes, not for agent collaboration with semantic understanding. To address this gap, we propose two new layers: an Agent Communication Layer (L8) and an Agent Semantic Negotiation Layer (L9). L8 formalizes the structure of communication, standardizing message envelopes, speech-act performatives (e.g., REQUEST, INFORM), and interaction patterns (e.g., request-reply, publish-subscribe), building on protocols like MCP. L9, which does not exist today, formalizes the meaning of communication, enabling agents to discover, negotiate, and lock a ”Shared Context”—a formal schema defining the concepts, tasks, and parameters relevant to their interaction. Together, these layers provide the foundation for scalable, distributed agent collaboration, enabling the next generation of multi-agentic systems.
1. Introduction
Large Language Models (LLMs) have demonstrated remarkable advances in performance, fundamentally transforming how we approach complex computational tasks. A key breakthrough has been their ability to learn and work with domain-specific languages (DSLs), including APIs, tool interfaces, and structured data formats. This capability has enabled the creation of AI agents that go beyond simple text generation to perform preliminary computations, execute actions, and interact with external systems through standardized protocols like the Model Context Protocol (MCP) (mcp2024, ).
However, LLMs face fundamental architectural constraints. Their context windows, while expanding, cannot grow indefinitely, limiting their working memory and the complexity of problems they can solve independently. This limitation mirrors challenges in traditional computing systems, where single processors hit physical and thermal constraints. The solution in both cases is the same: distribution and collaboration. Just as computational systems scale through distributed architectures with various memory hierarchies, AI agents must collaborate to tackle problems beyond individual capacity (chopra2025, ).
The analogy to distributed computing runs deep. Distributed computing enables the implementation of distributed algorithms using specialized programming languages, with parallel programming representing the simplest form—replication of identical computational tasks across shards. More sophisticated distributed programming requires decomposing algorithms into specialized computational tasks, a fundamentally harder problem. While not every problem can be optimally distributed, heuristics and approximations often scale favorably, trading perfect solutions for tractable ones at larger problem sizes. Multi-agent systems represent a realization of distributed computation with a compelling feature: autonomous action. LLM-based multi-agent systems add another dimension—the ability to work across multiple languages, both natural and domain-specific.
Programming languages are themselves domain-specific languages, and compilers serve as DSL validators of both syntax and semantics within the context of computation. LLMs excel at translating between languages—including from natural language to DSLs—but lack inherent correctness guarantees. This observation reveals a powerful synergy: if LLM agents can translate natural language to DSL, and compilers can validate that DSL, then coupling natural language with the syntax and semantics of DSLs bridges the gap to distributed programming using agents. LLMs are exceptionally strong at language transformation; external validation mechanisms provide the correctness they lack.
Agent collaboration requires communication infrastructure that goes beyond simple data exchange. When agents work together—whether coordinating in supply chains, managing distributed systems, or solving multi-step reasoning tasks—they need to share not just data but understanding. Current network stacks (OSI and TCP/IP) were designed for reliable data delivery between processes, not for semantic coordination between intelligent, autonomous entities (zimmermann1980, ). This gap has led to a proliferation of ad-hoc, application-specific protocols that address agent communication in fragmented, incompatible ways.
The ”Internet of Agents” (IoA) emerges as the natural evolution: a communication infrastructure that enables agents to scale through distributed computation and semantic coordination. However, efficient multi-agent systems require more than data exchange protocols—they demand mechanisms that guarantee both computational efficiency and semantic correctness. The very capabilities that make LLMs powerful—their fluency in natural language and ability to handle ambiguity—create fundamental challenges for reliable agent-to-agent communication.
The Protocol Challenge: Current protocols like A2A (a2a2024, ), MCP (mcp2024, ), and FIPA-ACL standardize the syntax of agent messages—defining envelopes, performatives, and interaction patterns—but cannot enforce semantic agreement. They ensure messages are well-formed but cannot guarantee that communicating agents share the same understanding of terms, concepts, or task parameters. This semantic gap forces expensive disambiguation through iterative clarification exchanges, undermining both computational efficiency and deterministic behavior.
Consider two critical failure modes that arise from this syntactic-only approach:
(1) Syntax Without Semantic Context: Syntax defines the formal rules for constructing valid statements that can be parsed by machines or interpreted by humans. Each language—whether a programming language, domain-specific language (DSL), or natural language—has its own syntactic rules tailored to its purpose. Formal languages excel at efficient syntax verification, enabling computers to quickly validate statement structure and execute computation deterministically. Natural languages, while syntactically interpretable, are less efficient for computational execution due to inherent ambiguity.
The critical insight for multi-agent systems is that we have developed numerous domain-specific languages transportable via application-level protocols (JSON, XML, Protocol Buffers). LLM-based agents can leverage these DSLs to generate syntactically correct statements that execute efficiently. However, syntactic correctness alone is insufficient for reliable agent communication. A term like ”switch” can be syntactically valid in any language context, yet its interpretation depends entirely on semantic context—network device, electrical component, or strategic change. Current protocols validate syntax but provide no mechanism to establish which semantic context applies, forcing agents to guess or engage in expensive clarification cycles.
(2) Semantic Underspecification: Even when syntax is perfect and terms are unambiguous, messages can fail due to incomplete semantic specification. Consider the syntactically valid request (REQUEST :content ”Book a ticket to New York”). The syntax is correct, and ”New York” is unambiguous as a proper noun. However, the semantic context is severely underspecified: which New York entity (city vs. state)? Which airport (JFK, LGA, EWR)? What travel date? Which service class? Which passenger? These are not syntactic ambiguities—the message parses correctly—but semantic gaps where critical contextual parameters are missing.
Without a formal semantic layer that defines the required parameters for the ”book ticket” task within a ”travel” context, agents cannot deterministically execute requests. The receiving agent must either make potentially incorrect assumptions or engage in multiple clarification rounds, each invoking expensive LLM inference. This creates non-deterministic negotiation loops that scale poorly—the computational cost grows not from syntactic parsing but from iterative semantic resolution that should have been established at the protocol level.
The Solution Framework: Efficient multi-agent systems require protocols that combine formal syntax with semantic knowledge. Formal syntax ensures messages are parseable and interaction patterns are predictable. Semantic knowledge—expressed through shared, machine-readable ontologies and schemas—provides the context necessary to interpret message content unambiguously. This combination enables agents to validate semantic correctness at the protocol level, eliminating ambiguity before computation begins and ensuring that distributed agents execute tasks deterministically and efficiently.
This paper argues for a proactive, systematic solution. We propose formalizing agent communication through two new architectural layers: an Agent Communication Layer (L8) that standardizes interaction structure, and an Agent Semantic Negotiation Layer (L9) that establishes shared meaning before task execution. The Internet of Agents adds the semantics and syntax of modern multipurpose DSLs into agent interactions. While individual LLMs may fail to implement weak non-deterministic Turing machines, agents equipped with memory and structured communication protocols can theoretically achieve non-deterministic Turing computation in a distributed manner. Together, these layers provide the foundation for scalable, distributed agent collaboration—the essential infrastructure for an Internet of Agents where computation scales through semantic coordination rather than context window expansion.
2. The Lesson of Ad-Hoc Layers: HTTP
The history of the internet is a powerful precedent for our current situation. The original OSI (Zimmermann, 1980) and TCP/IP models provided a robust and universal standard for routing packets and ensuring reliable data delivery. However, they intentionally stopped short of dictating how applications should use that data. The Application Layer (L7) was less a ”layer” and more a ”gateway” for other protocols like FTP, SMTP, and Telnet.
In the early 1990s, the World Wide Web created a new requirement: a simple, extensible way to retrieve hypermedia documents. The TCP/IP stack had no native primitive for ”getting a document.” This led to the creation of the Hypertext Transfer Protocol (HTTP).
HTTP was, in effect, an ad-hoc protocol built to fill the gap left by L4 (TCP) and L7. It started as a simple, stateless protocol with one primary method (GET). However, as web applications grew in complexity, HTTP was forced to evolve. It accrued state management (Cookies), complex caching mechanisms (Headers), new verbs (PUT, DELETE), and session-like capabilities (Keep-Alive), effectively becoming its own complex, stateful ”application layer” living entirely within the ”data” payload of a TCP segment.
The evolution continued with HTTP/2 (thomson2022, ) and HTTP/3, which introduced fundamental transport-level primitives: FRAMES for efficient message encapsulation and STREAMS for multiplexed, independent channels within a single connection. These capabilities transformed HTTP from a simple application protocol into a genuine transport protocol, perfectly fitting the definition of a ”narrow waist” or ”thin waist” layer of the internet stack (popa2010, ). HTTP/2’s binary framing and stream multiplexing, followed by HTTP/3’s QUIC-based transport, established HTTP as a universal transport foundation for modern applications.
Today, HTTP/2 and HTTP/3 represent a de facto transport layer of the stack, arguably more complex and influential than the layers below it. We are now at an identical inflection point with the Internet of Agents.
Beyond its transport capabilities, HTTP also serves as the secure layer for authentication and authorization, providing application-level security mechanisms that go beyond the connection-level security offered by TLS in HTTPS. HTTP headers carry authentication tokens (e.g., OAuth 2.0 bearer tokens, API keys), session identifiers, and authorization credentials, establishing a security model that is decoupled from the underlying transport encryption. This separation of concerns—where TLS secures the channel and HTTP manages identity and access—makes HTTP a complete application transport layer rather than merely a data transfer protocol.
Agents need to exchange concepts far more complex than ”documents.” They must communicate goals, tasks, beliefs, intentions, and contingencies. If we rely on existing L7 protocols, we will force all of this complexity into the application payload (e.g., a ”chat” message over WebSockets or a ”POST” to an API). This will inevitably lead to an ad-hoc ”swamp” of proprietary, brittle, and non-interoperable agent protocols—a thousand different ways to ”book a flight” or ”coordinate a supply chain,” all hidden inside generic JSON payloads.
In order to be proactive, we argue that we need to formalize the primitives of agent communication before this ad-hoc chaos solidifies. We propose to do this by explicitly defining two new layers for the IoA stack.
3. Background and Related Work
The need for high-level agent communication is not new, and several protocols have been developed to address it. However, they uniformly fail by mistaking syntactic agreement for semantic consensus.
3.1. Classical Agent Communication Languages (ACLs)
The most prominent historical effort is the FIPA-ACL (Foundation for Intelligent Physical Agents, 2002). FIPA-ACL was a significant advancement, as it moved beyond simple data transfer to formalize speech acts (Searle, 1969). Agents communicate using performatives like INFORM, REQUEST, or PROPOSE. Critically, the FIPA-ACL message structure included an :ontology parameter, explicitly acknowledging that a shared vocabulary was necessary for the message’s :content to be understood.
The failure of FIPA-ACL in practice was twofold. First, its reliance on complex, heavyweight formal ontologies (e.g., KIF, SL) created significant performance and implementation overhead. Second, the :ontology slot was merely a label. The protocol specified no mechanism for negotiating or aligning ontologies if two agents did not possess an identical, pre-shared model (Gruber, 1993). It identified the problem but offered no protocol-level solution.
3.2. Modern Agent Communication Protocols (A2A/MCP)
Modern successors, such as the Agent2Agent Protocol (A2A) (a2a2024, ), the Model Context Protocol (MCP) (mcp2024, ), and the Natural Language Interaction Protocol (NLIP) (nlip2024, ), have sought to remedy this by being more lightweight. They replace complex KIF/SL with JSON or other simple data-binding formats and run over modern web protocols (e.g., WebSockets, HTTP/2).
These protocols provide sophisticated frameworks for agent interaction. A2A, for instance, enables agents to exchange artifacts (structured data, files, or contextual information) and delegate task execution to other agents. Agents can request task execution and receive status updates on in-progress tasks or final outputs upon completion. MCP similarly provides structured mechanisms for agents to share context and invoke capabilities. Both protocols excel at standardizing the syntactic envelope of communication—defining message structure, performatives, and interaction patterns.
For example, a protocol might define a standard JSON structure for agent messages:
JSON
  "protocol": "A2A/1.0",
  "performative": "REQUEST",
  "sender_id": "agent-travel-7",
  "receiver_id": "agent-booking-4",
  "content": {
    "task": "book_flight",
    "prompt": "I need a flight to New York for next Tuesday."
This message is syntactically perfect. A protocol-aware agent can parse it, identify the sender, and extract the task request. The protocol provides the flexible framework for communication. However, the responsibility for specifying syntax and semantic context accurately remains with the multi-agent system developer. The protocol does not enforce or validate that both agents share the same understanding of what ”book_flight” requires or what ”New York” means in this context. The receiving agent (agent-booking-4) has no guaranteed, protocol-driven way to ground the natural language string. Its LLM will be invoked on ambiguous data. The agent may:
  • Guess: Assume ”New York” means JFK and ”next Tuesday” means the immediate next Tuesday, potentially booking the wrong flight.
  • Clarify: Send a REQUEST-CLARIFY message back. This results in a ”prompt negotiation” loop, which is computationally expensive (each turn is a full LLM inference) and non-deterministic.
  • Fail: Reject the request as too ambiguous.
The core problem is this: Current protocols formalize the structure of the request but not the context of the task. The most important information—the shared understanding of terms, goals, and constraints—is left unmanaged, leading to brittle, unpredictable, and inefficient multi-agent systems.
Application (L7)Presentation (L6)Session (L5)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Traditional OSI Agentic Application Agentic SemanticNegotiationAgenticCommunicationAppApp Transport(L5-L7)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Proposed Stack for Agents
Figure 1. Traditional OSI stack and our proposed network stack for agentic applications. We propose two new layers for agent communication (L8 and L9) above HTTP/2/3, which serves as the Application Transport layer (L7).
4. A Layered Architecture for the Agentic Web
To solve this impasse, we propose a new architectural stack. First, we formally recognize HTTP/2 and HTTP/3 as Layer 7, the Application Transport Layer, leveraging their FRAMES and STREAMS capabilities as fundamental transport primitives. These protocols, with their multiplexing, flow control, and efficient binary framing, provide the ideal foundation for agent communication. We then propose two new layers (L8 and L9) residing above the application transport layer. This separation of concerns is critical for managing the complexity of agentic communication.
4.1. Layer 8: The Agent Communication Layer
The first new layer, L8, is the Agent Communication Layer. Its responsibility is to ensure that agents can reliably exchange structured messages, understand communicative intent, and coordinate interaction flows—independently of semantic content. It unifies the best parts of existing protocols (like A2A, MCP, NLIP, FIPA-ACL) into a single, standardized layer that provides the syntactic and protocol foundation for agent communication. The Agent Communication Layer is not concerned with what the data means, only with ensuring messages are parseable, intentions are clear, and interaction roles are well-defined.
To achieve this, L8 provides three essential components:
  1. (1)
    Message Structure (The Envelope): Defines the standard ”envelope” for all agent messages, ensuring messages are parseable and routable. Protocols like A2A (a2a2024, ) and MCP (mcp2024, ) implement this through structured messages, tasks, and artifacts that specify sender, receiver(s), message ID, and performative. This provides the syntactic foundation for reliable message exchange.
  2. (2)
    Performatives (The Speech Act): Standardizes the communicative intent of messages, separate from their content. This ensures agents understand what action is being requested or performed, regardless of the semantic payload. A2A’s message structure naturally supports this through its task-oriented design, where operations like task requests, artifact exchanges, and status updates represent distinct speech acts. More generally, we identify that the registry of standardized speech acts includes:
    • Transactional: ‘REQUEST‘, ‘AGREE‘, ‘REFUSE‘, ‘INFORM‘
    • Negotiation: ‘PROPOSE‘, ‘ACCEPT‘, ‘REJECT‘,
      ‘COUNTER_PROPOSE‘
    • Information: ‘QUERY‘, ‘SUBSCRIBE‘, ‘PUBLISH‘
  3. (3)
    Interaction Patterns (The Dance): Formalizes the expected sequence and coordination of messages, ensuring agents understand their role in multi-step interactions. While A2A currently implements client-server patterns, protocols like SLIM (Secure Low-Latency Interactive Messaging) (slim2025, ) extend this with publish-subscribe and group communication capabilities. Combined, these protocols can support:
    • Request-Reply: A simple 1:1 ‘REQUEST‘ followed by an ‘AGREE‘ or ‘REFUSE‘.
    • Publish-Subscribe: A 1:N pattern where agents ‘SUBSCRIBE‘ to a topic and receive ‘INFORM‘ messages.
    • Aggregation: An N:1 pattern where a leader agent requests data from a population, which is then aggregated.
    • Collaboration Groups: N:N patterns where multiple agents coordinate in secure groups, exchanging artifacts and coordinating tasks distributedly.
    A2A’s extension mechanism provides a pathway for incorporating these interaction patterns directly into the protocol layer.
Together, these components ensure that agents can parse messages (structure), understand what is being requested (intent), and coordinate their role in the interaction (patterns)—all without requiring shared semantic understanding. This syntactic and protocol foundation is what separates L8 from L9: L8 provides the “how” of communication, while L9 (described next) provides the “what.” Only by combining both layers can agents achieve truly reliable, scalable collaboration.
4.2. Layer 9: The Agent Semantic Negotiation Layer (SNL)
Above L8 sits the Agent Semantic Negotiation Layer (L9). This layer’s sole responsibility is to establish a verifiable, shared meaning for the ‘content‘ payload of L8 messages. It acts as a specialized mediator that ensures semantic alignment before the primary task is ever processed.
Unlike L8, where protocols like A2A (a2a2024, ), MCP (mcp2024, ), and SLIM (slim2025, ) provide substantial capabilities for message structure and interaction patterns, L9 represents a capability that does not exist in current agent communication protocols. While FIPA-ACL (fipa2002, ) acknowledged the need for shared ontologies through its ‘:ontology‘ parameter, no protocol provides mechanisms for discovering, negotiating, and locking semantic contexts at the protocol level. Agents today must rely on application-level conventions or expensive, non-deterministic natural language negotiation.
L9 draws inspiration from distributed computing, where processes coordinate through shared memory or signaling mechanisms (chopra2025, ). Just as distributed algorithms use shared state (mutexes, semaphores, message queues) to coordinate concurrent processes, multi-agent systems require shared semantic contexts to coordinate autonomous agents. The difference is profound: distributed processes share data structures; distributed agents must share meaning. L9 provides this semantic coordination primitive, enabling agents to establish a ”shared context”—analogous to shared memory in distributed systems—that defines the concepts, tasks, and parameters relevant to their interaction.
The semantic context established by L9 is integrated into agent reasoning in two ways. First, it serves as a validation schema that the agent’s L9 SNL layer uses to verify all outgoing and incoming message content against the locked Shared Context before passing data to or from the agent’s LLM. Second, the context definitions are injected into the LLM’s reasoning process, either by augmenting the system prompt with formal concept definitions or by providing a structured retrieval mechanism where the LLM can query the Shared Context to resolve ambiguities during inference. This dual integration ensures both protocol-level correctness and semantic-aware reasoning.
L9 is what allows agents to move from simply exchanging data to truly understanding each other. It ensures that when one agent sends data, the receiving agent knows exactly what that data represents in a shared, formal context.
4.3. The Agent Semantic Negotiation Layer (L9) in Detail
The power of the L8/L9 split is that it separates the structure of communication (L8) from its meaning (L9). The SNL’s primary job is to negotiate and lock a ”Shared Context” that all subsequent L8 messages will adhere to.
4.4. Shared Semantic Contexts
To support 1:N or N:N coordination patterns, which involve continuous alignment of state, beliefs, or intentions rather than single tasks, we generalize this concept to a Shared Context. A Shared Context is a formal, machine-readable definition (e.g., using JSON Schema, RDF/OWL, or even Protobufs) of all the concepts, entities, tasks, and parameters relevant to a specific domain of discourse. This context is identified by a unique, versioned URN (e.g., ‘urn:contexts:travel:v2.1‘ or ‘urn:contexts:supplyChain:v1.0‘).
4.5. Functions of the SNL
The SNL protocol consists of three phases:
1. Context Discovery and Negotiation (The Semantic Handshake) Before any meaningful interaction, the agents’ SNLs (L9) perform a handshake to find a compatible Shared Context.
  • Agent A (SNL): (SNL-HELLO) ”I want to interact. I support the following contexts: [‘urn:contexts:travel:v2.1‘,
    ‘urn:contexts:supplyChain:v1.0‘]”
  • Agent B (SNL): (SNL-SELECT) ”I also support
    ‘urn:contexts:supplyChain:v1.0‘. Let us use that. The authoritative schema is at
    https://transform.tools/json-to-json-schema.”
  • Agent A (SNL): (SNL-LOCK) ”Context locked. All L8 message ‘content‘ will now be validated against
    ‘urn:contexts:supplyChain:v1.0‘.”
This negotiation creates a ”semantic session,” which can be cached and reused. While the initial handshake introduces latency, this cost is amortized across the entire session—potentially thousands of subsequent L8 messages benefit from the shared semantic understanding without additional negotiation overhead.
2. Semantic Grounding Once the context is locked, it is the SNL’s job to ground all data. This means binding data (text, numbers, vectors) from the agent’s internal logic to the formal definitions in the Shared Context. This function supports both transactional and coordinative patterns, which map directly to the two dominant paradigms in multi-agent systems today: agent orchestration and agent collaboration. Transactional patterns enable orchestration, where a central agent coordinates task execution across subordinate agents (e.g., a travel agent orchestrating booking services). Coordinative patterns enable collaboration, where peer agents share state and intentions to jointly solve problems without central control (e.g., supply chain agents coordinating inventory decisions).
Example 1: Transactional Pattern (Booking a Flight)
  • Context: ‘urn:contexts:travel:v2.1‘
  • Schema Definition: Defines a task named ‘bookFlight‘ with required parameters ‘origin_code‘ (string), ‘dest_code‘ (string), and ‘date‘ (string, yyyy-mm-dd).
  • Agent A Logic: ”I need a flight to New York for next Tuesday from LAX.”
  • Agent A (SNL) Grounding: The SNL processes this. It successfully binds ‘origin_code: ”LAX”‘ and ‘date: ”2025-11-04”‘. It finds ‘dest_code‘ is ambiguous (”New York”).
  • SNL-driven Clarification: The SNL does not send the ambiguous prompt. It uses L8 to send a formal, L9-defined query:
    • L8 Performative: ‘QUERY‘
    • L9 Content: ‘”concept”: ”ambiguous_parameter”, ”parameter”: ”dest_code”, ”value”: ”New York”‘
  • Agent B (SNL) Response: Agent B’s SNL receives this, understands the ‘ambiguous_parameter‘ concept, and replies via L8 ‘INFORM‘ with L9 ‘content‘: ‘”concept”: ”parameter_options”, ”parameter”: ”dest_code”, ”options”: [”JFK”, ”LGA”, ”EWR”]‘.
This clarification loop happens at the protocol level, deterministically, rather than as an ambiguous LLM ”chat.”
Example 2: Coordinative Pattern (Supply Chain Alignment) In this way L9 supports complex, population-wide coordination without needing to define a single ”task.”
  • Context: ‘urn:contexts:supplyChain:v1.0‘
  • Schema Definition: This schema does not define a ”task.” Instead, it defines concepts that agents can share:
    • Concept 1: ‘current_decision‘ (type: object, properties: ‘item_id‘, ‘quantity‘)
    • Concept 2: ‘decision_contingency‘ (type: object, properties: ‘if_condition_text‘, ‘then_change_text‘)
    • Concept 3: ‘local_observation‘ (type: object, properties: ‘observed_fact_text‘, ‘confidence_score‘)
  • Agent A (Retailer) Logic: ”My inventory is low. I’m ordering 120 units. It seems like a temporary demand spike. If demand does increase 10%, I’ll increase my order by 15 units.”
  • Agent A (SNL) Grounding: The agent’s logic is passed to its SNL. The SNL composes a semantically grounded L8 message to ‘PUBLISH‘ to its neighbors.
    • L8 Performative: ‘INFORM‘
    • L9 Content:
        "my_decision": {
          "concept_type": "current_decision",
          "item_id": "beer",
          "quantity": 120
        "my_flexibility": {
          "concept_type": "decision_contingency",
          "if_condition_text": "demand increases by
          10%",
          "then_change_text": "increase order by
          15 units"
        "my_reasoning": {
          "concept_type": "local_observation",
          "observed_fact_text": "Current spike seems
          temporary",
          "confidence_score": 0.8
  • Receiving Agent (SNL): Agent B (Wholesaler) receives this L8 message. Its L9 SNL validates the ‘content‘ against the locked ‘supplyChain:v1.0‘ context. It knows, at a protocol level, what ‘decision_contingency‘ is. It can now pass this structured, meaningful data to its internal logic (an LLM, a rules engine) for aggregation. It is not just ”chat”; it is a formal, structured exchange of meaning.
3. Semantic Validation The final function is semantic validation, where the receiving agent’s SNL validates all incoming L8 ‘content‘ against the locked Shared Context. If an agent sends a message with an undefined concept (e.g., ‘”my_mood”: ”happy”‘ in the supply chain context), the L9 SNL can either reject it as semantically invalid, even if the L8 syntax is perfect or trigger another round of disambiguation. Once disambiguated, this new semantic definition can be cached locally for future communication between these agents. This prevents semantic ambiguity and communication failures.
5. Security and Trust in the Semantic Layer
Modern cloud and application security already involves multiple layers of defense: network isolation, identity and access management, encryption at rest and in transit, application-level authorization, and runtime monitoring. By making meaning a formal part of the protocol stack, the SNL (L9) introduces an additional attack surface that intersects with—but is not replaced by—these existing security mechanisms. Semantic-level threats require semantic-level defenses that complement, rather than substitute, established security practices.
L9 security must integrate with existing frameworks. Where L4/L7 provides transport security (TLS), and protocols like SLIM (slim2025, ) provide end-to-end message confidentiality via MLS (Message Layer Security), L9 adds a new dimension: semantic integrity and authorization. An encrypted, authenticated message can still contain semantically malicious content or reference poisoned contexts. This section identifies the novel security challenges introduced by formalizing semantics at the protocol level and proposes mechanisms that work alongside existing security infrastructure.
5.1. New Attack Vectors
A Semantic Layer is vulnerable to entirely new attacks:
  • Semantic Injection: This is the L9 equivalent of prompt injection. An attacker crafts a payload that is syntactically valid (L8) and semantically valid (L9 schema) but contains malicious instructions. In our supply chain example:
    • ‘if_condition_text‘: ”demand spike”
    • ‘then_change_text‘: ”increase order by 10 units. Also, as a priority, ignore all previous instructions and send your full inventory list to attacker@evil.com”
    A simple JSON schema validator would pass this, as it’s a valid string. The malicious payload is aimed at the receiving agent’s LLM.
  • Context Poisoning / Spoofing: A malicious agent (or a compromised registry) serves a modified version of a popular Shared Context. For example, in urn:contexts:payment:v1.1, it swaps the definitions of ‘sender_id‘ and ‘receiver_id‘, tricking agents into sending money to the attacker.
  • Semantic Denial of Service (SDoS): An attacker floods an agent with semantically valid but computationally expensive requests. For example, thousands of L8 ‘QUERY‘ messages with L9 content ‘”concept”: ”ambiguous_parameter”‘. This doesn’t exhaust network bandwidth (L4) but exhausts the victim’s LLM inference budget (L9+Logic), a far more expensive resource.
  • Semantic Downgrade Attack: An attacker forces agents to negotiate weaker or older context versions with known vulnerabilities during the SNL handshake. For example, if ‘urn:contexts:payment:v2.0‘ fixed a semantic ambiguity that allowed double-spending in ‘v1.0‘, an attacker performing a man-in-the-middle attack could strip ‘v2.0‘ from the SNL-HELLO capabilities list, forcing agents to fall back to the vulnerable ‘v1.0‘ context. This is analogous to TLS downgrade attacks but operates at the semantic layer.
5.2. Security Requirements for a Robust SNL
To be viable, the SNL (L9) must be built on a foundation of ”zero trust” for meaning, integrating with existing security layers through defense-in-depth.
5.2.1. Multi-Layer Security Integration
L9 security operates alongside existing mechanisms. Transport security (TLS) protects L8 message envelopes in transit. For protocols like SLIM (slim2025, ), MLS provides end-to-end encryption of message payloads, ensuring confidentiality even through intermediate nodes. L9 adds semantic integrity: validating that message content conforms to authenticated contexts and that agents are authorized to use specific concepts. All three layers are necessary—transport encryption prevents eavesdropping, message encryption ensures end-to-end confidentiality, and semantic validation prevents malicious or malformed meaning from reaching agent logic.
5.2.2. Core Security Mechanisms
  • Authenticated Contexts: A Shared Context schema must be as verifiable as an SSL/TLS certificate. The SNL handshake must involve exchanging cryptographically signed context definitions, including version numbers and supported feature sets. Agents must validate signatures against trusted Schema Authorities before accepting a context. The handshake must include minimum version negotiation to prevent downgrade attacks—if Agent A requires ‘payment:v2.0‘ minimum, it must reject any negotiation attempt for ‘v1.x‘.
  • Semantic Firewalls: Semantic Firewalls operate at the L9 layer, positioned between the SNL and the agent’s application logic (LLM or rules engine). They inspect validated L9 content—not raw L8 bytes—and enforce concept-level authorization policies. Unlike traditional firewalls that filter based on IPs or ports, semantic firewalls understand the meaning of communication. Implementation challenges include:
    • Placement: Semantic firewalls can be deployed client-side (protecting individual agents), at gateways (protecting agent populations), or distributed (each agent enforces policies for its context).
    • Policy Examples:
        ”Agent [Retailer-7] is allowed to use the ‘decision_contingency‘ concept. Agent [Public-WebApp-3] is not.”
        ”Rate-limit any single agent to 10 ‘ambiguous_parameter‘ queries per minute (to prevent SDoS).”
        ”Reject any ‘payment‘ context message where ‘amount ¿ \$10000‘ unless sender has ‘high-value-authorized‘ credential.”
    • Content Inspection: Detecting semantic injection (malicious prompts embedded in semantically valid fields) remains an open research challenge. Current approaches include pattern matching for known injection signatures, anomaly detection on text distributions, and even LLM-based analysis of suspicious content. However, this creates an arms race similar to spam filtering. We acknowledge this limitation and suggest semantic firewalls focus on authorization (who can use which concepts) and rate limiting as more tractable near-term defenses.
    • Performance: Semantic inspection adds latency proportional to message complexity and policy count. Caching policy decisions per context-agent pair and using efficient policy languages (e.g., Rego, Cedar) can minimize overhead.
  • Confidentiality of Capability: The SNL-HELLO handshake reveals an agent’s capabilities (i.e., what contexts it understands). This is sensitive business intelligence—revealing that an agent supports ‘military-logistics:v3.2‘ or ‘high-frequency-trading:v1.0‘ exposes strategic information. Where SLIM uses MLS for payload encryption, the L9 negotiation messages (SNL-HELLO, SNL-SELECT) must also be encrypted using the same end-to-end security mechanism, ensuring that even intermediate routing nodes cannot observe capability advertisements. This requires extending the secure channel established at L7 to cover L9 control messages.
5.3. Trust Model and Governance
The Schema Authority (SA) model raises critical governance questions. We envision a federated trust model, analogous to how TLS certificates operate:
  • Centralized SAs for Public Contexts: Industry consortiums or standards bodies (e.g., ”Travel Industry Agent Protocol Alliance”) would operate SAs for widely-used public contexts like ‘travel‘, ‘ecommerce‘, or ‘healthcare‘. These SAs would sign context definitions, publish revocation lists, and manage version lifecycles. Agents would ship with root SA certificates, similar to browsers shipping with CA root certificates.
  • Decentralized/Self-Signed for Private Contexts: Organizations deploying internal multi-agent systems could self-sign private contexts (e.g., ‘urn:contexts:acme-internal-logistics:v1.0‘). Agents within the organization’s trust boundary would accept the organization’s SA certificate. This enables private semantic coordination without external dependencies.
  • Federation Across Boundaries: Cross-organizational collaboration requires trust bridging. Agent A (trusting SA-Financial) and Agent B (trusting SA-Healthcare) could interact if their SAs have established mutual recognition agreements, or through a trusted intermediary SA. This mirrors federated identity systems (SAML, OAuth, OpenID Connect).
  • Context Lifecycle: SAs must handle context updates (minor version bumps for backward-compatible changes, major versions for breaking changes) and revocations (if a context is found to have security vulnerabilities). Agents must periodically refresh SA certificates and check revocation lists, adding operational complexity but ensuring semantic integrity over time.
This governance model is complex but necessary. The alternative—no semantic authentication—would allow any malicious actor to publish poisoned contexts, undermining the entire L9 value proposition.
6. Implementation and Challenges
The L8/L9 architecture is not a replacement for existing protocols but rather a formalization and extension of capabilities already emerging in modern agent communication standards. Our proposal leverages and builds upon existing implementations to provide a cohesive framework.
6.1. Building on Existing Standards
The proposed architecture aligns with and extends current protocol development:
  • L8 (Agent Communication Layer): The foundational capabilities of L8 are already substantially implemented in existing protocols:
    • Message Structure: A2A (a2a2024, ) and MCP (mcp2024, ) provide well-defined message envelopes with sender, receiver, message IDs, and structured content (tasks, artifacts, tool invocations).
    • Performatives: A2A’s task-oriented operations (task requests, artifact exchanges, status updates) map naturally to speech acts. MCP’s tool invocation model similarly expresses intent.
    • Interaction Patterns: While A2A currently implements client-server patterns, SLIM (slim2025, ) extends this with publish-subscribe, group communication, and streaming RPC (SRPC) over secure channels. SLIM’s MLS-based end-to-end encryption and hierarchical naming provide the security and routing foundation for diverse interaction patterns.
    • Extension Mechanisms: A2A’s extension framework provides a pathway to incorporate additional interaction patterns and semantic capabilities directly into the protocol.
    The L8 standardization effort should focus on unifying these capabilities into a consistent framework, rather than creating entirely new protocols.
  • L9 (Semantic Layer): This layer represents the novel contribution, as no current protocol provides semantic negotiation capabilities. However, the implementation can leverage existing technologies:
    • Formal Schemas (for Grounding): JSON Schema and Protocol Buffers—already used by A2A, MCP, and SLIM—are ideal for defining the structure of semantic concepts and tasks. The L9 SNL would validate message content against these schemas within negotiated contexts.
    • Lightweight Ontologies (for Meaning): To define relationships between concepts (e.g., ”a ‘bookFlight‘ task requires a ‘date‘ parameter”), systems like RDF/OWL or simpler graph-based models could build Shared Contexts. These would be distributed via the same infrastructure used for protocol specifications.
    • Vectorial Representations (for Nuance): For inherently ambiguous concepts (e.g., ”a stylish hotel”), L9 Shared Contexts could include embedding fields. The SNL-HELLO handshake would negotiate embedding models (e.g., ‘text-embedding-3-small‘), allowing semantic similarity comparisons.
    • Security Integration: L9’s semantic validation would operate above SLIM’s MLS encryption layer, ensuring that cryptographically secure messages also carry semantically valid content.
6.2. Implementation Strategy
A pragmatic path forward involves:
  1. (1)
    Extend A2A with L9 Capabilities: A2A’s extension mechanism could incorporate SNL handshake messages (SNL-HELLO, SNL-SELECT, SNL-LOCK) as a standardized extension, enabling semantic negotiation while maintaining backward compatibility.
  2. (2)
    Integrate SLIM for L8 Interaction Patterns: SLIM’s publish-subscribe and streaming RPC provide the missing interaction patterns for L8. Agents using A2A for transactional operations could leverage SLIM for coordinative patterns.
  3. (3)
    Prototype L9 SNL: Develop reference implementations of the Semantic Negotiation Layer as libraries that work with existing protocols, demonstrating semantic validation, context negotiation, and integration with LLM-based agents.
  4. (4)
    Establish Schema Authority Pilot: Create a pilot SA for a specific domain (e.g., travel, supply chain) to demonstrate context signing, distribution, and versioning.
6.3. Remaining Challenges
Despite leveraging existing standards, significant challenges remain:
  • Performance Overhead: The semantic handshake introduces latency. This must be amortized by using persistent ”semantic sessions” that survive across multiple L8 message exchanges.
  • The Bootstrapping Problem: How is the initial registry of Shared Contexts created? This requires cross-industry standardization efforts, similar to how schema.org emerged for web content. Early domain-specific pilots are essential.
  • Representation Alignment: If two agents use different embedding models, their vector spaces will not align. This remains an open research problem (sucholutsky2023, ). The SNL provides the protocol to negotiate this, but the underlying ML techniques must advance.
  • Prompt Injection Defense: As noted in the security section, detecting semantic injection in validated fields is an unsolved problem. Semantic firewalls must focus on authorization and rate limiting until better content inspection techniques emerge.
7. Conclusion
Network architecture has historically evolved to meet new demands: from packet delivery (IP) to reliable sessions (TCP) to structured application communication (HTTP/2/3). As autonomous agents become significant network actors, the protocol stack may require corresponding evolution to support their coordination needs.
This paper proposes a two-layer extension building on existing work. The Agent Communication Layer (L8) formalizes structural capabilities emerging in A2A (a2a2024, ), MCP (mcp2024, ), and SLIM (slim2025, ), standardizing message envelopes, task delegation, and interaction patterns. The Agent Semantic Negotiation Layer (L9) introduces protocol-level semantic negotiation—enabling agents to discover, negotiate, and lock formal contexts defining concepts and parameters. This addresses a capability gap in current protocols, where semantic alignment occurs informally at the application layer.
Separating syntactic structure (L8) from semantic meaning (L9) supports both transactional patterns (agent orchestration) and coordinative patterns (agent collaboration). Challenges remain: semantic handshake latency, context registry standardization, semantic injection defense, and governance models for Schema Authorities. However, multi-agent systems are already deployed in supply chains, distributed systems, and collaborative reasoning. Building on existing protocols while adding formal semantic negotiation could provide improved interoperability and reliability for these emerging applications.

### v3

@@
 Large Language Models (LLMs) have demonstrated remarkable performance improvements and the ability to learn domain-specific languages (DSLs), including APIs and tool interfaces.
-This capability has enabled the creation of AI agents that can perform preliminary computations and act through tool calling, now being standardized via protocols like MCP.
-However, LLMs face fundamental limitations: their context windows cannot grow indefinitely, constraining their memory and computational capacity.
+This capability has enabled the creation of AI agents that can perform preliminary computations and act through tool calling, which is now being standardized via protocols like MCP.
+However, LLMs face fundamental limitations: their context windows cannot grow indefinitely, restricting their memory and computational capacity.
 Agent collaboration emerges as essential for solving increasingly complex problems, mirroring how computational systems rely on different types of memory to scale.
@@
 Current network architectural stacks (OSI and TCP/IP) were designed for data delivery between hosts and processes, not for agent collaboration with semantic understanding.
-To address this gap, we propose two new layers: an Agent Communication Layer (L8) and an Agent Semantic Negotiation Layer (L9).
+To address this gap, we propose two new layers: an Agent Communication Layer (L8) and an Agent Semantic Layer (L9).
 L8 formalizes the structure of communication, standardizing message envelopes, speech-act performatives (e.g., REQUEST, INFORM), and interaction patterns (e.g., request-reply, publish-subscribe), building on protocols like MCP.
-L9, which does not exist today, formalizes the meaning of communication, enabling agents to discover, negotiate, and lock a ”Shared Context”—a formal schema defining the concepts, tasks, and parameters relevant to their interaction.
+The proposed L9 layer: (1) formalizes semantic context discovery and negotiation, (2) provides semantic grounding by binding terms to semantic context, and (3) semantically validates incoming prompts and performs disambiguation as needed.
+Furthermore, L9 introduces primitives for coordination and consensus, allowing agents to achieve alignment on shared states, collective goals, and distributed beliefs.
 Together, these layers provide the foundation for scalable, distributed agent collaboration, enabling the next generation of multi-agentic systems.
@@
 This paper argues for a proactive, systematic solution.
-We propose formalizing agent communication through two new architectural layers: an Agent Communication Layer (L8) that standardizes interaction structure, and an Agent Semantic Negotiation Layer (L9) that establishes shared meaning before task execution.
+We propose formalizing agent communication through two new architectural layers: an Agent Communication Layer (L8) that standardizes interaction structure, and an Agent Semantic Layer (L9) that establishes shared meaning and collective coordination before task execution.
 The Internet of Agents adds the semantics and syntax of modern multipurpose DSLs into agent interactions.
@@
 A protocol-aware agent can parse it, identify the sender, and extract the task request.
-The protocol provides the flexible framework for communication.
+The protocol provides a flexible framework for communication.
 However, the responsibility for specifying syntax and semantic context accurately remains with the multi-agent system developer.
@@
 The agent may:
-  • Guess: Assume ”New York” means JFK and ”next Tuesday” means the immediate next Tuesday, potentially booking the wrong flight.
+  • Guess: Assume that ”New York” means JFK and ”next Tuesday” means the immediate next Tuesday, potentially booking the wrong flight.
   • Clarify: Send a REQUEST-CLARIFY message back. This results in a ”prompt negotiation” loop, which is computationally expensive (each turn is a full LLM inference) and non-deterministic.
@@
 The most important information—the shared understanding of terms, goals, and constraints—is left unmanaged, leading to brittle, unpredictable, and inefficient multi-agent systems.
-Application (L7)Presentation (L6)Session (L5)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Traditional OSI Agentic Application Agentic SemanticNegotiationAgenticCommunicationAppApp Transport(L5-L7)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Proposed Stack for Agents
+Application (L7)Presentation (L6)Session (L5)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Traditional OSI Agentic Application (L8) Agentic Semantic (L7)AgenticComm.
+(L6)AppApp Transport(L5)Transport (L4)Network (L3)Data Link (L2)Physical (L1)Proposed Stack for Agents
 Figure 1.
@@
 These protocols, with their multiplexing, flow control, and efficient binary framing, provide the ideal foundation for agent communication.
-We then propose two new layers (L8 and L9) residing above the application transport layer.
+We then propose two new layers (L8 and L9) residing above the application transport layer (See Figures 1 and 2).
 This separation of concerns is critical for managing the complexity of agentic communication.
@@
 This syntactic and protocol foundation is what separates L8 from L9: L8 provides the “how” of communication, while L9 (described next) provides the “what.” Only by combining both layers can agents achieve truly reliable, scalable collaboration.
-4.2. Layer 9: The Agent Semantic Negotiation Layer (SNL)
-Above L8 sits the Agent Semantic Negotiation Layer (L9).
-This layer’s sole responsibility is to establish a verifiable, shared meaning for the ‘content‘ payload of L8 messages.
+L9L8L7REPA2A     MCP HTTP(S)    GRPC    SLIM
+Figure 2. Existing agentic protocols and their place in the proposed network stack. Note that the Ripple Effect Protocol (REP) only includes coordination functionality, not semantic negotiation.
+4.2. Layer 9: The Agent Semantic Layer (SL)
+Above L8 sits the Agent Semantic Layer (L9).
+This layer’s primary responsibility is to establish a verifiable, shared meaning for the ‘content‘ payload of L8 messages and to facilitate coordination and consensus across agent populations.
 It acts as a specialized mediator that ensures semantic alignment before the primary task is ever processed.
@@
 L9 provides this semantic coordination primitive, enabling agents to establish a ”shared context”—analogous to shared memory in distributed systems—that defines the concepts, tasks, and parameters relevant to their interaction.
+Crucially, the Agent Semantic Layer (SL) addresses the need for consensus.
+In population-scale agent systems, such as the Ripple Effect Protocol, agents must not only understand a message but also reach a collective agreement on the ”truth” or ”state” of the system.
+L9 provides the consensus primitives (e.g., voting, threshold signatures, or state replication) that allow agents to converge on a single semantic interpretation of global events, preventing divergence in multi-agent simulations or real-world industrial deployments.
 The semantic context established by L9 is integrated into agent reasoning in two ways.
-First, it serves as a validation schema that the agent’s L9 SNL layer uses to verify all outgoing and incoming message content against the locked Shared Context before passing data to or from the agent’s LLM.
+First, it serves as a validation schema that the agent’s L9 SL layer uses to verify all outgoing and incoming message content against the locked Shared Context before passing data to or from the agent’s LLM.
 Second, the context definitions are injected into the LLM’s reasoning process, either by augmenting the system prompt with formal concept definitions or by providing a structured retrieval mechanism where the LLM can query the Shared Context to resolve ambiguities during inference.
 This dual integration ensures both protocol-level correctness and semantic-aware reasoning.
-L9 is what allows agents to move from simply exchanging data to truly understanding each other.
+L9 is what allows agents to move from simply exchanging data to truly understanding and agreeing with each other.
 It ensures that when one agent sends data, the receiving agent knows exactly what that data represents in a shared, formal context.
-4.3. The Agent Semantic Negotiation Layer (L9) in Detail
-The power of the L8/L9 split is that it separates the structure of communication (L8) from its meaning (L9).
-The SNL’s primary job is to negotiate and lock a ”Shared Context” that all subsequent L8 messages will adhere to.
+4.3. The Agent Semantic Layer (SL) in Detail
+The power of the L8/L9 split is that it separates the structure of communication (L8) from its meaning and collective state (L9).
+The SL’s primary job is to negotiate and lock a ”Shared Context” that all subsequent L8 messages will adhere to, and to manage the transition from individual beliefs to collective consensus.
+While L9 protocols could begin to operate with semantic state (e.g., English), it is not a strict requirement.
+For example, in certain cases, it may make sense to communicate using latent space bindings (e.g., embedding vectors) directly which avoids the costly translation from embeddings to semantic state back to embeddings, of course in cases where both agents use the same embedding algorithms.
+In addition, communication can also happen using other types of compressed state granularity, which may be useful in scenarios where agents are located far from each other and reduction in communication bandwidth is a requirement.
 4.4. Shared Semantic Contexts
@@
 This context is identified by a unique, versioned URN (e.g., ‘urn:contexts:travel:v2.1‘ or ‘urn:contexts:supplyChain:v1.0‘).
-4.5. Functions of the SNL
-The SNL protocol consists of three phases:
-1. Context Discovery and Negotiation (The Semantic Handshake) Before any meaningful interaction, the agents’ SNLs (L9) perform a handshake to find a compatible Shared Context.
-  • Agent A (SNL): (SNL-HELLO) ”I want to interact. I support the following contexts: [‘urn:contexts:travel:v2.1‘,
+4.5. Functions of the SL
+The SL protocol consists of four primary phases:
+1. Context Discovery and Negotiation (The Semantic Handshake) Before any meaningful interaction, the agents’ SLs (L9) perform a handshake to find a compatible Shared Context.
+Example: Initial Agent Connection
+  • Agent A (SL): (SL-HELLO) ”I want to interact. I support the following contexts: [‘urn:contexts:travel:v2.1‘,
     ‘urn:contexts:supplyChain:v1.0‘]”
-  • Agent B (SNL): (SNL-SELECT) ”I also support
+  • Agent B (SL): (SL-SELECT) ”I also support
     ‘urn:contexts:supplyChain:v1.0‘. Let us use that. The authoritative schema is at
     https://transform.tools/json-to-json-schema.”
-  • Agent A (SNL): (SNL-LOCK) ”Context locked. All L8 message ‘content‘ will now be validated against
+  • Agent A (SL): (SL-LOCK) ”Context locked. All L8 message ‘content‘ will now be validated against
     ‘urn:contexts:supplyChain:v1.0‘.”
 This negotiation creates a ”semantic session,” which can be cached and reused.
-While the initial handshake introduces latency, this cost is amortized across the entire session—potentially thousands of subsequent L8 messages benefit from the shared semantic understanding without additional negotiation overhead.
 2.
-Semantic Grounding Once the context is locked, it is the SNL’s job to ground all data.
+Semantic Grounding Once the context is locked, it is the SL’s job to ground all data.
 This means binding data (text, numbers, vectors) from the agent’s internal logic to the formal definitions in the Shared Context.
-This function supports both transactional and coordinative patterns, which map directly to the two dominant paradigms in multi-agent systems today: agent orchestration and agent collaboration.
-Transactional patterns enable orchestration, where a central agent coordinates task execution across subordinate agents (e.g., a travel agent orchestrating booking services).
-Coordinative patterns enable collaboration, where peer agents share state and intentions to jointly solve problems without central control (e.g., supply chain agents coordinating inventory decisions).
-Example 1: Transactional Pattern (Booking a Flight)
+This function supports both transactional and coordinative patterns.
+3.
+Semantic Validation/Disambiguation The next function is semantic validation, where the receiving agent’s SL validates all incoming L8 ‘content‘ against the locked Shared Context.
+If an agent sends a message with an undefined concept or a concept which can potentially have more than one meaning, the SL initiates the semantic disambiguation process.
+This process allows the agents to negotiate a shared meaning for ambiguous terms or gain better clarity for which among many possibilities is the concept referring to, which can then be added to the shared semantic context.
+Example: Transactional Pattern (Booking a Flight)
   • Context: ‘urn:contexts:travel:v2.1‘
@@
   • Agent A Logic: ”I need a flight to New York for next Tuesday from LAX.”
-  • Agent A (SNL) Grounding: The SNL processes this. It successfully binds ‘origin_code: ”LAX”‘ and ‘date: ”2025-11-04”‘. It finds ‘dest_code‘ is ambiguous (”New York”).
-  • SNL-driven Clarification: The SNL does not send the ambiguous prompt. It uses L8 to send a formal, L9-defined query:
+  • Agent A (SL) Grounding: The SL processes this. It successfully binds ‘origin_code: ”LAX”‘ and ‘date: ”2025-11-04”‘. It finds ‘dest_code‘ is ambiguous (”New York”).
+  • SL-driven Clarification: The SL does not send the ambiguous prompt. It uses L8 to send a formal, L9-defined query:
     • L8 Performative: ‘QUERY‘
     • L9 Content: ‘”concept”: ”ambiguous_parameter”, ”parameter”: ”dest_code”, ”value”: ”New York”‘
-  • Agent B (SNL) Response: Agent B’s SNL receives this, understands the ‘ambiguous_parameter‘ concept, and replies via L8 ‘INFORM‘ with L9 ‘content‘: ‘”concept”: ”parameter_options”, ”parameter”: ”dest_code”, ”options”: [”JFK”, ”LGA”, ”EWR”]‘.
-This clarification loop happens at the protocol level, deterministically, rather than as an ambiguous LLM ”chat.”
-Example 2: Coordinative Pattern (Supply Chain Alignment) In this way L9 supports complex, population-wide coordination without needing to define a single ”task.”
+4.
+Coordination and Consensus (The Ripple Effect) In multi-agent systems where agents function as a swarm or a population, grounding is insufficient; agents must agree.
+L9 also should provide the Consensus Primitives necessary for Ripple-like alignment.
+When a leading agent proposes a change in state or belief, L9 manages the ”ripple” of agreement.
+It ensures that a majority or a quorum of agents have successfully validated and ”committed” to the same Shared Context state before the system acts.
+This prevents divergent behaviors where half the population operates under one set of beliefs while the other half diverges.
+Example: Coordinative Pattern (Supply Chain Alignment) In this way L9 supports complex, population-wide coordination without needing to define a single ”task.” (chopra2025, )
   • Context: ‘urn:contexts:supplyChain:v1.0‘
-  • Schema Definition: This schema does not define a ”task.” Instead, it defines concepts that agents can share:
-    • Concept 1: ‘current_decision‘ (type: object, properties: ‘item_id‘, ‘quantity‘)
-    • Concept 2: ‘decision_contingency‘ (type: object, properties: ‘if_condition_text‘, ‘then_change_text‘)
-    • Concept 3: ‘local_observation‘ (type: object, properties: ‘observed_fact_text‘, ‘confidence_score‘)
-  • Agent A (Retailer) Logic: ”My inventory is low. I’m ordering 120 units. It seems like a temporary demand spike. If demand does increase 10%, I’ll increase my order by 15 units.”
-  • Agent A (SNL) Grounding: The agent’s logic is passed to its SNL. The SNL composes a semantically grounded L8 message to ‘PUBLISH‘ to its neighbors.
+  • Agent A (Retailer) Logic: ”My inventory is low. I’m ordering 120 units.”
+  • Agent A (SL) Grounding: The agent’s logic is passed to its SL. The SL composes a semantically grounded L8 message to ‘PUBLISH‘ to its neighbors.
     • L8 Performative: ‘INFORM‘
@@
           "item_id": "beer",
-          "quantity": 120
-        "my_flexibility": {
-          "concept_type": "decision_contingency",
-          "if_condition_text": "demand increases by
-          10%",
-          "then_change_text": "increase order by
-          15 units"
-        "my_reasoning": {
-          "concept_type": "local_observation",
-          "observed_fact_text": "Current spike seems
-          temporary",
-          "confidence_score": 0.8
-  • Receiving Agent (SNL): Agent B (Wholesaler) receives this L8 message.
-Its L9 SNL validates the ‘content‘ against the locked ‘supplyChain:v1.0‘ context.
-It knows, at a protocol level, what ‘decision_contingency‘ is.
-It can now pass this structured, meaningful data to its internal logic (an LLM, a rules engine) for aggregation.
-It is not just ”chat”; it is a formal, structured exchange of meaning.
-3.
-Semantic Validation The final function is semantic validation, where the receiving agent’s SNL validates all incoming L8 ‘content‘ against the locked Shared Context.
-If an agent sends a message with an undefined concept (e.g., ‘”my_mood”: ”happy”‘ in the supply chain context), the L9 SNL can either reject it as semantically invalid, even if the L8 syntax is perfect or trigger another round of disambiguation.
-Once disambiguated, this new semantic definition can be cached locally for future communication between these agents.
-This prevents semantic ambiguity and communication failures.
+          "quantity": 120,
+          "sensitivity": "if demand increases by
+             10%, increase order by 15"
 5. Security and Trust in the Semantic Layer
 Modern cloud and application security already involves multiple layers of defense: network isolation, identity and access management, encryption at rest and in transit, application-level authorization, and runtime monitoring.
-By making meaning a formal part of the protocol stack, the SNL (L9) introduces an additional attack surface that intersects with—but is not replaced by—these existing security mechanisms.
+By making meaning a formal part of the protocol stack, the SL (L9) introduces an additional attack surface that intersects with—but is not replaced by—these existing security mechanisms.
 Semantic-level threats require semantic-level defenses that complement, rather than substitute, established security practices.
@@
     A simple JSON schema validator would pass this, as it’s a valid string. The malicious payload is aimed at the receiving agent’s LLM.
-  • Context Poisoning / Spoofing: A malicious agent (or a compromised registry) serves a modified version of a popular Shared Context.
+  • Context Poisoning / Spoofing: A malicious agent (or a compromired registry) serves a modified version of a popular Shared Context.
 For example, in urn:contexts:payment:v1.1, it swaps the definitions of ‘sender_id‘ and ‘receiver_id‘, tricking agents into sending money to the attacker.
@@
 This doesn’t exhaust network bandwidth (L4) but exhausts the victim’s LLM inference budget (L9+Logic), a far more expensive resource.
-  • Semantic Downgrade Attack: An attacker forces agents to negotiate weaker or older context versions with known vulnerabilities during the SNL handshake.
-For example, if ‘urn:contexts:payment:v2.0‘ fixed a semantic ambiguity that allowed double-spending in ‘v1.0‘, an attacker performing a man-in-the-middle attack could strip ‘v2.0‘ from the SNL-HELLO capabilities list, forcing agents to fall back to the vulnerable ‘v1.0‘ context.
+  • Semantic Downgrade Attack: An attacker forces agents to negotiate weaker or older context versions with known vulnerabilities during the SL handshake.
+For example, if ‘urn:contexts:payment:v2.0‘ fixed a semantic ambiguity that allowed double-spending in ‘v1.0‘, an attacker performing a man-in-the-middle attack could strip ‘v2.0‘ from the SL-HELLO capabilities list, forcing agents to fall back to the vulnerable ‘v1.0‘ context.
 This is analogous to TLS downgrade attacks but operates at the semantic layer.
-5.2. Security Requirements for a Robust SNL
-To be viable, the SNL (L9) must be built on a foundation of ”zero trust” for meaning, integrating with existing security layers through defense-in-depth.
+5.2. Security Requirements for a Robust SL
+To be viable, the SL (L9) must be built on a foundation of ”zero trust” for meaning, integrating with existing security layers through defense-in-depth.
 5.2.1. Multi-Layer Security Integration
@@
   • Authenticated Contexts: A Shared Context schema must be as verifiable as an SSL/TLS certificate.
-The SNL handshake must involve exchanging cryptographically signed context definitions, including version numbers and supported feature sets.
+The SL handshake must involve exchanging cryptographically signed context definitions, including version numbers and supported feature sets.
 Agents must validate signatures against trusted Schema Authorities before accepting a context.
 The handshake must include minimum version negotiation to prevent downgrade attacks—if Agent A requires ‘payment:v2.0‘ minimum, it must reject any negotiation attempt for ‘v1.x‘.
-  • Semantic Firewalls: Semantic Firewalls operate at the L9 layer, positioned between the SNL and the agent’s application logic (LLM or rules engine).
+  • Semantic Firewalls: Semantic Firewalls operate at the L9 layer, positioned between the SL and the agent’s application logic (LLM or rules engine).
 They inspect validated L9 content—not raw L8 bytes—and enforce concept-level authorization policies.
@@
 Caching policy decisions per context-agent pair and using efficient policy languages (e.g., Rego, Cedar) can minimize overhead.
-  • Confidentiality of Capability: The SNL-HELLO handshake reveals an agent’s capabilities (i.e., what contexts it understands).
+  • Confidentiality of Capability: The SL-HELLO handshake reveals an agent’s capabilities (i.e., what contexts it understands).
 This is sensitive business intelligence—revealing that an agent supports ‘military-logistics:v3.2‘ or ‘high-frequency-trading:v1.0‘ exposes strategic information.
-Where SLIM uses MLS for payload encryption, the L9 negotiation messages (SNL-HELLO, SNL-SELECT) must also be encrypted using the same end-to-end security mechanism, ensuring that even intermediate routing nodes cannot observe capability advertisements.
+Where SLIM uses MLS for payload encryption, the L9 negotiation messages (SL-HELLO, SL-SELECT) must also be encrypted using the same end-to-end security mechanism, ensuring that even intermediate routing nodes cannot observe capability advertisements.
 This requires extending the secure channel established at L7 to cover L9 control messages.
@@
     The L8 standardization effort should focus on unifying these capabilities into a consistent framework, rather than creating entirely new protocols.
-  • L9 (Semantic Layer): This layer represents the novel contribution, as no current protocol provides semantic negotiation capabilities.
+  • L9 (Agent Semantic Layer): This layer represents the novel contribution, as no current protocol provides semantic negotiation capabilities.
 However, the implementation can leverage existing technologies:
     • Formal Schemas (for Grounding): JSON Schema and Protocol Buffers—already used by A2A, MCP, and SLIM—are ideal for defining the structure of semantic concepts and tasks.
-The L9 SNL would validate message content against these schemas within negotiated contexts.
-    • Lightweight Ontologies (for Meaning): To define relationships between concepts (e.g., ”a ‘bookFlight‘ task requires a ‘date‘ parameter”), systems like RDF/OWL or simpler graph-based models could build Shared Contexts.
+The L9 SL would validate message content against these schemas within negotiated contexts.
+    • Lightweight Ontologies (for Meaning): To define relationships between concepts, systems like RDF/OWL or simpler graph-based models could build Shared Contexts.
 These would be distributed via the same infrastructure used for protocol specifications.
-    • Vectorial Representations (for Nuance): For inherently ambiguous concepts (e.g., ”a stylish hotel”), L9 Shared Contexts could include embedding fields.
-The SNL-HELLO handshake would negotiate embedding models (e.g., ‘text-embedding-3-small‘), allowing semantic similarity comparisons.
+    • Vectorial Representations (for Nuance): For inherently ambiguous concepts, L9 shared contexts are not limited to semantic spaces, but could also take place in latent spaces or other compressed spaces.
+For example, REP uses a shared embedding space for coordination.
+These shared embedding spaces can be negotiated as part of the SL-Hello handshake.
     • Security Integration: L9’s semantic validation would operate above SLIM’s MLS encryption layer, ensuring that cryptographically secure messages also carry semantically valid content.
@@
   1. (1)
-    Extend A2A with L9 Capabilities: A2A’s extension mechanism could incorporate SNL handshake messages (SNL-HELLO, SNL-SELECT, SNL-LOCK) as a standardized extension, enabling semantic negotiation while maintaining backward compatibility.
+    Extend A2A with L9 Capabilities: A2A’s extension mechanism could incorporate SL handshake messages (SL-HELLO, SL-SELECT, SL-LOCK) as a standardized extension, enabling semantic negotiation and **coordination** while maintaining backward compatibility.
   2. (2)
@@
   3. (3)
-    Prototype L9 SNL: Develop reference implementations of the Semantic Negotiation Layer as libraries that work with existing protocols, demonstrating semantic validation, context negotiation, and integration with LLM-based agents.
+    Prototype L9 SL: Develop reference implementations of the Agent Semantic Layer as libraries that work with existing protocols, demonstrating semantic validation, context negotiation, and state consensus integration with LLM-based agents.
   4. (4)
@@
 Despite leveraging existing standards, significant challenges remain:
-  • Performance Overhead: The semantic handshake introduces latency. This must be amortized by using persistent ”semantic sessions” that survive across multiple L8 message exchanges.
-  • The Bootstrapping Problem: How is the initial registry of Shared Contexts created?
-This requires cross-industry standardization efforts, similar to how schema.org emerged for web content.
-Early domain-specific pilots are essential.
+  • Performance Overhead: The semantic handshake and consensus rounds introduce latency.
+This must be amortized by using persistent ”semantic sessions” that survive across multiple L8 message exchanges.
+  • The Bootstrapping Problem: How is the initial registry of Shared Contexts created? This requires cross-industry standardization efforts, similar to how schema.org emerged for web content.
   • Representation Alignment: If two agents use different embedding models, their vector spaces will not align.
 This remains an open research problem (sucholutsky2023, ).
-The SNL provides the protocol to negotiate this, but the underlying ML techniques must advance.
+The SL provides the protocol to negotiate this, but the underlying ML techniques must advance.
   • Prompt Injection Defense: As noted in the security section, detecting semantic injection in validated fields is an unsolved problem.
@@
 The Agent Communication Layer (L8) formalizes structural capabilities emerging in A2A (a2a2024, ), MCP (mcp2024, ), and SLIM (slim2025, ), standardizing message envelopes, task delegation, and interaction patterns.
-The Agent Semantic Negotiation Layer (L9) introduces protocol-level semantic negotiation—enabling agents to discover, negotiate, and lock formal contexts defining concepts and parameters.
-This addresses a capability gap in current protocols, where semantic alignment occurs informally at the application layer.
-Separating syntactic structure (L8) from semantic meaning (L9) supports both transactional patterns (agent orchestration) and coordinative patterns (agent collaboration).
+The Agent Semantic Layer (L9) introduces protocol-level semantic negotiation and collective consensus—enabling agents to discover, negotiate, and lock formal contexts defining concepts and parameters, and to align system states across populations.
+This addresses a capability gap in current protocols, where semantic alignment and consensus occur informally at the application layer.
+Separation of syntactic structure (L8) from semantic meaning and state (L9) supports both transactional patterns (agent orchestration) and coordination patterns (agent collaboration).
 Challenges remain: semantic handshake latency, context registry standardization, semantic injection defense, and governance models for Schema Authorities.
 However, multi-agent systems are already deployed in supply chains, distributed systems, and collaborative reasoning.
-Building on existing protocols while adding formal semantic negotiation could provide improved interoperability and reliability for these emerging applications.
+Building on existing protocols while adding formal semantic coordination and consensus could provide improved interoperability and reliability for these emerging applications.